OpenStudy (anonymous):

I like really need help! you have six $1 bills, two $10 bills, and four $20 bills in your wallet. You select a bill at random. Without replacing the bill, you choose a second bill. what is P($1, then $10)? (A)77/190 (B)3/100 (C)3/95 (D)2/5

OpenStudy (anonymous):

You should construct a probability tree in order to answer your quest :)

OpenStudy (anonymous):

How many total bills do you have? How many of these are $1 bills? \[P($1)=\frac{\text{number of }$1}{\text{total number of bills}}\] How many bills are left when you remove the 1$ bill? How many of these $10 bills? \[P($1,\text{ then }$10)=P($1)\times\frac{\text{number of }$10}{\text{total number of bills}-1}\]